OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 400x8 – 36.

OpenStudy (anonymous):

First, look for any factors in common between the terms.

OpenStudy (anonymous):

36 is not quite divisible by 8.

OpenStudy (anonymous):

Those are from 36, and 400 doesn't have any multiples of 3, so that leaves....

OpenStudy (anonymous):

No, 4. We're looking for common factors.

OpenStudy (anonymous):

400X^8-36= 4{ 100X^8 - 9}} = 4{ ( 10X^4)2 - 3^2} = 4( 10X^4+3)( 10X^4-3)

OpenStudy (anonymous):